# Quantum Sieving for Code-Based Cryptanalysis and Its Limitations for ISD

### v1

@@ 4.1 Probability $p$ of Forming a Solution Pair
|$\frac{|\{(\mathbf{x},\mathbf{y})\in(\mathcal{S}_{w}^{n})^{2}\colon|\mathbf{x}\land\mathbf{c}|=\alpha,|\mathbf{y}\land\mathbf{c}|=\alpha,|\mathbf{x}\land\mathbf{y}|=w/2\}|}{|\{(\mathbf{x},\mathbf{y})\in(\mathcal{S}_{w}^{n})^{2}\colon|\mathbf{x}\land\mathbf{y}|=w/2\}|}=\frac{\sum_{e}{v\choose\alpha}{n-v\choose w-\alpha}{\alpha\choose e}{v-\alpha\choose\alpha-e}{w-\alpha\choose w/2-e}{n-(w+v-\alpha)\choose w/2-\alpha+e}}{{n\choose w}{w\choose w/2}{n-w\choose w/2}}$
where $e$ ranges over all possible values of $|\mathbf{x}\land\mathbf{y}\land\mathbf{c}|$. Now, using that ($\frac{{v\choose\alpha}{n-v\choose w-\alpha}}{{n\choose w}}=\frac{{w\choose\alpha}{n-w\choose v-\alpha}}{{n\choose v}}$, (${w\choose\alpha}{\alpha\choose e}{w-\alpha\choose w/2-e}={w\choose w/2}{w/2\choose e}{w-w/2\choose\alpha-e}$, and (${n-w\choose v-\alpha}{v-\alpha\choose\alpha-e}{(n-w)-(v-\alpha)\choose w/2-(\alpha-e)}={n-w\choose w/2}{w/2\choose\alpha-e}{(n-w)-w/2\choose(v-\alpha)-(\alpha-e)}$, it can be seen that this indeed equals $\frac{|\mathcal{W}_{v,\alpha}^{n}(\mathbf{x}^{\prime},\mathbf{y}^{\prime})|}{{n\choose v}}$. ∎
@@ 4.2 Quantum Algorithm for FindSolutions Using Grover’s Algorithm
The statement follows from repeatedly applying Grover’s algorithm (Theorem 2.1) with the set of all (${B\choose 2}$ pairs as the search space. The expected number of solutions is $t:={B\choose 2}p$. Using a coupon-collector argument, it follows that all solutions can be found in expected time $\tilde{O}(\sqrt{t{B\choose 2}})=\tilde{O}(B^{2}\sqrt{p})$. The algorithm classically stores the $B$ list vectors. Grover’s algorithm requires quantum access to them (QRACM) and a number of qubits that is polynomial in $n$. ∎
@@ 4.3 Quantum Algorithm for FindSolutions Using Quantum Walks
We will now replace the role of Grover’s algorithm with quantum-walk techniques. In particular, we describe a FindSolutions subroutine that searches for solution pairs ($\mathbf{x},\mathbf{y}$ such that $|\mathbf{x}+\mathbf{y}|=w$) inside a given list $\mathcal{B}_{\alpha}(\mathbf{c})$ using a quantum walk. To speed up the search for solution pairs, we add a layer of locality-sensitive filtering (LSF) as part of the data of each vertex. This idea was inspired by [CL21], and we show that we can indeed transfer most of their ideas from the Euclidean metric to the Hamming metric. We analyze the complexity of the resulting quantum walk, and consider a version of the walk where we apply the ‘sparsification’ technique from [CL21], resulting in a different balance of the complexity parameters.
@@
With a slight abuse of notation, we write $\mathbb{E}[|\mathcal{VC}_{\beta}(\mathbf{x})|])$ as shorthand for $\max(1,\mathbb{E}[|\mathcal{VC}_{\beta}(\mathbf{x})|])$ to simplify the expressions of the time and memory complexity. We do the same for $\mathbb{E}[|\mathcal{B}_{\beta}(\mathbf{c}^{\prime})|]$.
@@ 4.4 Adding a Layer of Filtering
Let $\mathcal{B}_{\alpha}(\mathbf{c})$ be an instance of Problem 4.1 with parameters $(B,n,w,v,\alpha)$. There exists an integer $e^{*}$ such that the expected number of pairs $(\mathbf{x},\mathbf{y})\in\mathcal{B}_{\alpha}(\mathbf{c})^{2}$ satisfying $|\mathbf{x}+\mathbf{y}|=w$ is asymptotically equal to the expected number of pairs $(\mathbf{x},\mathbf{y})\in\mathcal{B}_{\alpha}(\mathbf{c})^{2}$ satisfying $|\mathbf{x}+\mathbf{y}|=w$ and $|\mathbf{x}\land\mathbf{y}\land\mathbf{c}|=e^{*}$.
In other words, Problem 4.1 is asymptotically equivalent to the following problem:
@@
Let :=$\mathcal{L}^{\prime}:=\{\pi_{\mathbf{c}}(\mathbf{x})\colon\mathbf{x}\in\mathcal{B}_{\alpha}(\mathbf{c})\}$, where we view each $\pi_{\mathbf{c}}(\mathbf{x}):=\mathbf{x}\land\mathbf{c}$ as a vector in $S_{\alpha}^{v}$ (i.e., we ignore all coefficients $(\mathbf{x}\land\mathbf{c})_{i}$ where $\mathbf{c}_{i}=0$). Note that the vectors in $\mathcal{L}^{\prime}$ are independently and uniformly distributed over $S_{\alpha}^{v}$. In particular, the problem of finding all $\mathbf{x}^{\prime},\mathbf{y}^{\prime}\in\mathcal{L}^{\prime}$ satisfying $|\mathbf{x}^{\prime}\land\mathbf{y}^{\prime}|=e^{*}$ is a variant of an NNS problem (where the target weight $e^{*}$ does not necessarily equal the weight $\alpha$ of the vectors in the input list $\mathcal{L}^{\prime}$), so we can apply the LSF techniques from [DEEK24]. Furthermore, the set of those pairs $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$ (or actually the corresponding $(\mathbf{x},\mathbf{y})$) forms a superset of the set $\{(\mathbf{x},\mathbf{y})\in\mathcal{B}_{\alpha}(\mathbf{c})^{2}\colon|\mathbf{x}+\mathbf{y}|=w\}$. Therefore, to asymptotically solve Problem 4.1, it suffices to find all $\mathbf{x}^{\prime},\mathbf{y}^{\prime}\in\mathcal{L}^{\prime}$ satisfying $|\mathbf{x}^{\prime}\land\mathbf{y}^{\prime}|=e^{*}$, and keep those for which it also holds that $|\mathbf{x}+\mathbf{y}|=w$.
@@
where $p(e):=\Pr_{\mathbf{x},\mathbf{y}\in_{R}\mathcal{B}_{\alpha}(\mathbf{c})}[|\mathbf{x}+\mathbf{y}|=w\text{ and }|\mathbf{x}\land\mathbf{y}\land\mathbf{c}|=e]$. In particular, $p=\tilde{\Theta}(p(e^{*}))$, where $e^{*}:=\mathrm{argmax}_{e}p(e)$. [^10] The value of $e^{*}$ can be computed numerically. For instance, see [Car20, DEEK24]. (Note that an explicit formula for $p(e)$ is given in Lemma 4.1.) This implies that
$\mathbb{E}[|\{(\mathbf{x},\mathbf{y})\in\mathcal{B}_{\alpha}(\mathbf{c})^{2}\colon|\mathbf{x}+\mathbf{y}|=w\}|]\quad\text{ and }\quad\mathbb{E}[|\{(\mathbf{x},\mathbf{y})\in\mathcal{B}_{\alpha}(\mathbf{c})^{2}\colon|\mathbf{x}+\mathbf{y}|=w\text{ and }|\mathbf{x}\land\mathbf{y}\land\mathbf{c}|=e^{*}\}|]$
are asymptotically equal, as we wanted to show. ∎
@@ 4.6.2 Memory complexity analysis
We need to store the bucket $\mathcal{B}_{\alpha}(\mathbf{c})$ (or, equivalently, the list $\mathcal{L}^{\prime}$), so the algorithm uses classical space $\tilde{O}(|\mathcal{B}_{\alpha}(\mathbf{c})|)$.
@@
The quantum walk needs quantum access to the bucket $\mathcal{B}_{\alpha}(\mathbf{c})$, stored classically and of size $M_{QRACM}=\tilde{O}(|\mathcal{B}_{\alpha}(\mathbf{c})|)$.
@@
The quantum walk stores a superposition of vertices. Each vertex contains $s$ elements and some additional data, including the buckets of the second layer. (The other components of the data will not dominate the QRAQM cost, so we can safely ignore those.) Within each vertex, a vector $\mathbf{x}^{\prime}$ is inserted in $|\mathcal{VC}_{\beta}(\mathbf{x}^{\prime})|$ buckets. All together, we need to store this on a quantum register of size $M_{Q}=s+s\cdot\mathbb{E}[|\mathcal{VC}_{\beta}(\mathbf{x}^{\prime})|]$. This is the required number of qubits to run the quantum walk. Since we use the data structures from [Amb07] (or [BJLM13]) to efficiently delete and insert vectors in the quantum superposition, we require this whole quantum register to allow efficient QRAQM, hence we need $M_{QRAQM}=M_{Q}$. ∎
@@ 5 Numerical Results
Table 1: The exponents of the asymptotic runtime (i.e., $t$ s.t. runtime is $2^{tn+o(n)}$) and corresponding memory exponents for the hardest instances of NNS for codes (i.e., for those $\omega$ that yield the highest runtime for each considered algorithm), and the asymptotic runtime and memory of their lattice analogs.
@@
We also implemented a variant of QW + LSF + Sparsification where we add the reusable quantum-walk techniques from [BCSS23] (in a similar way as they apply their techniques to [CL21]). Similar to their result on lattice sieving, we only obtain a minor improvement, namely an exponent of $t=0.1169$ instead of $0.1171$. (In lattice sieving, [BCSS23] obtain an exponent of $t=0.2563$ instead of $0.2570$.) Our implementation is available in our repository. We leave the elaboration of the analytical details of this approach as potential future work.
@@
While our work only focused on improving the asymptotic time complexities of sieving algorithms, it is also important to quantify the memory costs required by these algorithms, and minimize them. In particular, note that the improved runtime obtained when moving from classical to Grover, and subsequently to the quantum-walk algorithms, comes at the cost of an increased memory complexity. In particular, the use of Grover requires quantum memory and QRACM access. The use of quantum walks additionally requires QRAQM, which is an even stronger assumption on the memory model. Note that this trade-off between improved runtime and stronger memory requirements is consistent with the observations from other quantum algorithms in code-based (and lattice-based) cryptanalysis.
@@ 6 On the Application to Quantum Information-Set Decoding
Information-set decoding (ISD) algorithms are known as the most efficient generic attacks against the decoding problem for a wide range of parameters. [^15] For a more specific range of parameters, there exist more efficient attacks such as statistical decoding [CDMT22].. Each of these can be seen as improvements of the algorithm originally introduced by Prange[Pra62]. We view ISD algorithms as a framework, as formulated in [FS09]. Specifically, following [GJN23] and [DEEK24] we see the sieving algorithm as a subroutine of an ISD algorithm within the framework. We refer to the resulting algorithm as SievingISD. In this section, we analyze a quantum analog of SievingISD.
@@ 6.1 ISD and SievingISD
2 while False do
3       Choose $\mathbf{x}\in\mathcal{S}_{n^{\prime}}^{n}$ uniformly at random and repeat if $\dim(\pi_{\mathbf{x}}(\mathcal{C}))\neq k$.
4       $\mathcal{L}\leftarrow\mathcal{A}(\pi_{\mathbf{x}}(\mathcal{C}),w^{\prime})$.
@@ 6.2 Quantum ISD and Quantum SievingISD
We will now discuss quantum analogs of the ISD framework and SievingISD. Quantum ISD algorithms (e.g., [KT17, Kir18]) are based on the idea that one can apply amplitude amplification (AA) [BHMT02] to speed up the search for the solution of the decoding problem over multiple iterations of the ISD. Furthermore, the subroutine $\mathcal{A}$ is allowed to be quantum. It results in the following quantum analog of [DEEK24, Theorem 3.1]. Here, and in the remainder of Section 6, we only focus on the runtime of the (quantum) ISD algorithms, not its space usage.
@@
Let $\mathcal{C}$ be a a unique decoding instance with parameters $(n,k,w)$. Let $n^{\prime}>k$ and let $w^{\prime}\leq n^{\prime}$. Suppose $\mathcal{A}$ is a algorithm that returns $N$ independent and uniformly random weight-$w^{\prime}$ codewords in a given $[n^{\prime},k]$ binary linear code, where $N\leq{n^{\prime}\choose w^{\prime}}/2^{n^{\prime}-k}$.

### v2

@@ 4.1 Probability $p$ of Forming a Solution Pair
|$\,\frac{|\{(\mathbf{x},\mathbf{y})\in(\mathcal{S}_{w}^{n})^{2}\colon|\mathbf{x}\land\mathbf{c}|=\alpha,|\mathbf{y}\land\mathbf{c}|=\alpha,|\mathbf{x}\land\mathbf{y}|=w/2\}|}{|\{(\mathbf{x},\mathbf{y})\in(\mathcal{S}_{w}^{n})^{2}\colon|\mathbf{x}\land\mathbf{y}|=w/2\}|}$
$=$ $\,\frac{\sum_{e}{v\choose\alpha}{n-v\choose w-\alpha}{\alpha\choose e}{v-\alpha\choose\alpha-e}{w-\alpha\choose w/2-e}{n-(w+v-\alpha)\choose w/2-\alpha+e}}{{n\choose w}{w\choose w/2}{n-w\choose w/2}}$
where $e$ ranges over all possible values of $|\mathbf{x}\land\mathbf{y}\land\mathbf{c}|$. Now, using the equalities ($\frac{{v\choose\alpha}{n-v\choose w-\alpha}}{{n\choose w}}=\frac{{w\choose\alpha}{n-w\choose v-\alpha}}{{n\choose v}}$, (${w\choose\alpha}{\alpha\choose e}{w-\alpha\choose w/2-e}={w\choose w/2}{w/2\choose e}{w-w/2\choose\alpha-e}$, and (${n-w\choose v-\alpha}{v-\alpha\choose\alpha-e}{(n-w)-(v-\alpha)\choose w/2-(\alpha-e)}={n-w\choose w/2}{w/2\choose\alpha-e}{(n-w)-w/2\choose(v-\alpha)-(\alpha-e)}$, it can be seen that this indeed equals $\frac{|\mathcal{W}_{v,\alpha}^{n}(\mathbf{x}^{\prime},\mathbf{y}^{\prime})|}{{n\choose v}}$. ∎
@@ 4.2 Quantum Algorithm for FindSolutions Using Grover’s Algorithm
The statement follows from repeatedly applying Grover’s algorithm (Theorem 2.1) with the set of all (${B\choose 2}$ pairs as the search space. The expected number of solutions is $t:={B\choose 2}p$. Using a coupon-collector argument, it follows that all solutions can be found in expected time $\tilde{O}(\sqrt{t{B\choose 2}})=\tilde{O}(B^{2}\sqrt{p})$. The $B$ list vectors are classically stored, and Grover’s algorithm uses quantum access to them (QRACM); it also uses a polynomial (in $n$) number of qubits. ∎
@@ 4.3 Quantum Algorithm for FindSolutions Using Quantum Walks
We will now replace Grover’s algorithm with quantum-walk techniques. In particular, we describe a FindSolutions subroutine that searches for solution pairs ($\mathbf{x},\mathbf{y}$ such that $|\mathbf{x}+\mathbf{y}|=w$) inside a given list $\mathcal{B}_{\alpha}(\mathbf{c})$ using a quantum walk. To speed up the search for solution pairs, we add a layer of locality-sensitive filtering (LSF) as part of the data of each vertex. This idea was inspired by [CL21], and we show that we can indeed transfer most of their ideas from the Euclidean metric to the Hamming metric. We analyze the complexity of the resulting quantum walk, and consider a version of the walk where we apply the ‘sparsification’ technique from [CL21], resulting in a different balance of the complexity parameters.
@@
With a slight abuse of notation, we write $\mathbb{E}[|\mathcal{VC}_{\beta}(\mathbf{x})|])$ for $\max(1,\mathbb{E}[|\mathcal{VC}_{\beta}(\mathbf{x})|])$ to simplify the expressions for time and memory complexity. We do the same for $\mathbb{E}[|\mathcal{B}_{\beta}(\mathbf{c}^{\prime})|]$.
@@ 4.4 Adding a Layer of Filtering
Let $\mathcal{B}_{\alpha}(\mathbf{c})$ be an instance of Problem 4.1 with parameters $(B,n,w,v,\alpha)$. There exists an integer $e^{*}$ such that the expected number of pairs $(\mathbf{x},\mathbf{y})\in\mathcal{B}_{\alpha}(\mathbf{c})^{2}$ satisfying $|\mathbf{x}+\mathbf{y}|=w$ is equal, up to a multiplicative factor that is linear in $n$, to the expected number of pairs $(\mathbf{x},\mathbf{y})\in\mathcal{B}_{\alpha}(\mathbf{c})^{2}$ satisfying $|\mathbf{x}+\mathbf{y}|=w$ and $|\mathbf{x}\land\mathbf{y}\land\mathbf{c}|=e^{*}$.
In other words, Problem 4.1 is essentially equivalent to the following problem:
@@
Let :=$\mathcal{L}^{\prime}:=\{\pi_{\mathbf{c}}(\mathbf{x})\colon\mathbf{x}\in\mathcal{B}_{\alpha}(\mathbf{c})\}$, where we view each $\pi_{\mathbf{c}}(\mathbf{x}):=\mathbf{x}\land\mathbf{c}$ as a vector in $S_{\alpha}^{v}$ (i.e., we ignore all coefficients $(\mathbf{x}\land\mathbf{c})_{i}$ where $\mathbf{c}_{i}=0$). Note that the vectors in $\mathcal{L}^{\prime}$ are independently and uniformly distributed over $S_{\alpha}^{v}$. That is, the problem of finding all $\mathbf{x}^{\prime},\mathbf{y}^{\prime}\in\mathcal{L}^{\prime}$ satisfying $|\mathbf{x}^{\prime}\land\mathbf{y}^{\prime}|=e^{*}$ is a variant of an NNS problem (where the target weight $e^{*}$ does not necessarily equal the weight $\alpha$ of the vectors in the input list $\mathcal{L}^{\prime}$), so we can apply the LSF techniques from [DEEK24]. Furthermore, the set of those pairs $(\mathbf{x}^{\prime},\mathbf{y}^{\prime})$ (or actually the corresponding $(\mathbf{x},\mathbf{y})$) forms a superset of the set $\{(\mathbf{x},\mathbf{y})\in\mathcal{B}_{\alpha}(\mathbf{c})^{2}\colon|\mathbf{x}+\mathbf{y}|=w\}$. Therefore, to asymptotically solve Problem 4.1, it suffices to find all $\mathbf{x}^{\prime},\mathbf{y}^{\prime}\in\mathcal{L}^{\prime}$ satisfying $|\mathbf{x}^{\prime}\land\mathbf{y}^{\prime}|=e^{*}$, and keep those for which it also holds that $|\mathbf{x}+\mathbf{y}|=w$.
@@
where $p(e):=\Pr_{\mathbf{x},\mathbf{y}\in_{R}\mathcal{B}_{\alpha}(\mathbf{c})}[|\mathbf{x}+\mathbf{y}|=w\text{ and }|\mathbf{x}\land\mathbf{y}\land\mathbf{c}|=e]$. In particular, it holds that $p(e^{*})\leq p\leq p(e^{*})\cdot n/2$, where $e^{*}:=\mathrm{argmax}_{e}p(e)$. [^10] The value of $e^{*}$ can be computed numerically. For instance, see [Car20, DEEK24]. (Note that an explicit formula for $p(e)$ is given in Lemma 4.1.) This implies that $\mathbb{E}[|\{(\mathbf{x},\mathbf{y})\in\mathcal{B}_{\alpha}(\mathbf{c})^{2}\colon|\mathbf{x}+\mathbf{y}|=w\}|]$ and $\mathbb{E}[|\{(\mathbf{x},\mathbf{y})\in\mathcal{B}_{\alpha}(\mathbf{c})^{2}\colon|\mathbf{x}+\mathbf{y}|=w\text{ and }|\mathbf{x}\land\mathbf{y}\land\mathbf{c}|=e^{*}\}|]$ are equal up to a multiplicative factor that is linear in $n$, as we wanted to show. ∎
@@ 4.6.2 Memory complexity analysis
The algorithm stores the bucket $\mathcal{B}_{\alpha}(\mathbf{c})$ (or, equivalently, the list $\mathcal{L}^{\prime}$), so the algorithm uses classical space $\tilde{O}(|\mathcal{B}_{\alpha}(\mathbf{c})|)$.
@@
The quantum walk uses quantum access to the bucket $\mathcal{B}_{\alpha}(\mathbf{c})$, stored classically and of size $M_{QRACM}=\tilde{O}(|\mathcal{B}_{\alpha}(\mathbf{c})|)$.
@@
The quantum walk stores a superposition of vertices. Each vertex contains $s$ elements and some additional data, including the buckets of the second layer. (The other components of the data will not dominate the QRAQM cost, so we can safely ignore those.) Within each vertex, a vector $\mathbf{x}^{\prime}$ is inserted in $|\mathcal{VC}_{\beta}(\mathbf{x}^{\prime})|$ buckets. Altogether, this is stored in a quantum register of size $M_{Q}=s+s\cdot\mathbb{E}[|\mathcal{VC}_{\beta}(\mathbf{x}^{\prime})|]$ (which is the number of qubits used). Since we use the data structures from [Amb07] (or [BJLM13]) to efficiently delete and insert vectors in superposition, we use QRAQM of size $M_{QRAQM}=M_{Q}$. ∎
@@ 5 Numerical Results
Table 1: The exponents of the asymptotic runtime (i.e., $t$ s.t. runtime is $2^{tn+o(n)}$) and corresponding memory exponents for the hardest instances of NNS for codes (i.e., for those $\omega$ that yield the highest runtime for each considered algorithm), and the asymptotic runtime and memory of their lattice analogs. A dash ‘-’ means that the memory type is not used.
@@
Our repository also includes code to estimate the complexity of a variant of QW+LSF+ Sparsification where we incorporate the reusable quantum-walk techniques from [BCSS23] (in a similar way as they apply their techniques to [CL21]). Similar to their result on lattice sieving, we only obtain a minor improvement, namely an exponent of $t=0.1169$ instead of $0.1171$. (In lattice sieving, [BCSS23] obtain an exponent of $t=0.2563$ instead of $0.2570$.) We leave the elaboration of the analytical details of this approach as potential future work.
@@
While our work only focused on improving the asymptotic time complexities of sieving algorithms, it is also important to quantify the memory used by these algorithms, and minimize it. In particular, note that the improved runtime obtained when moving from classical to Grover, and subsequently to the quantum-walk algorithms, comes at the cost of an increased memory complexity and extra assumption about the memory model. Specifically, the Grover version uses quantum memory and QRACM access; quantum walks additionally use QRAQM, which is an even stronger assumption about the memory model. Note that this trade-off between improved runtime and assumption regarding the memory model is consistent with the observations from other quantum algorithms in code-based (and lattice-based) cryptanalysis.
@@ 6 On the Application to Quantum Information-Set Decoding
Information-set decoding (ISD) algorithms are known as the most efficient generic attacks against the decoding problem for a wide range of parameters. [^15] For a more specific range of parameters, there exist more efficient attacks such as statistical decoding [CDMT22]. Each of these can be seen as improvements of the algorithm originally introduced by Prange[Pra62]. We view ISD algorithms as a framework, as formulated in [FS09]. Specifically, following [GJN23] and [DEEK24] we see the sieving algorithm as a subroutine of an ISD algorithm within the framework. We refer to the resulting algorithm as SievingISD. In this section, we analyze a quantum analog of SievingISD.
@@ 6.1 ISD and SievingISD
2 while True do
3       Choose $\mathbf{x}\in\mathcal{S}_{n^{\prime}}^{n}$ uniformly at random and repeat if $\dim(\pi_{\mathbf{x}}(\mathcal{C}))\neq k$
4       $\mathcal{L}\leftarrow\mathcal{A}(\pi_{\mathbf{x}}(\mathcal{C}),w^{\prime})$
@@ 6.2 Quantum ISD and Quantum SievingISD
We will now discuss quantum analogs of the ISD framework and SievingISD. Quantum ISD algorithms (e.g., [KT17, Kir18]) are based on the idea that one can apply amplitude amplification (AA) [BHMT02] to speed up the search for the solution of the decoding problem over multiple iterations of the ISD. Furthermore, the subroutine $\mathcal{A}$ is allowed to be quantum. It results in the following quantum analog of [DEEK24, Theorem 3.1]. Here, and in the remainder of Section 6, we only focus on the runtime of (quantum) ISD algorithms, not their space usage.
@@
Let $\mathcal{C}$ be a unique decoding instance with parameters $(n,k,w)$. Let $n^{\prime}>k$ and let $w^{\prime}\leq n^{\prime}$. Suppose $\mathcal{A}$ is an algorithm that returns $N$ independent and uniformly random weight-$w^{\prime}$ codewords in a given $[n^{\prime},k]$ binary linear code, where $N\leq{n^{\prime}\choose w^{\prime}}/2^{n^{\prime}-k}$.